# Robust LLM Alignment via Distributionally Robust Direct Preference Optimization

### v2

@@ 5.1 Estimation Error for WDPO
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$ and $K=\lvert\log\sigma(-4\beta B)\rvert$, $\lambda$ is the regularity number defined in Assumption 2.
@@
Strong duality of Wasserstein DRO (see Gao and Kleywegt, (2022) and Corollary 1) helps us reduce the difference $\left\lvert\mathcal{L}^{\mathrm{W}}(\theta;\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta;\rho)\right\rvert$ to the concentration $\lvert\mathbb{E}_{z\sim\mathsf{P}^{o}}[l_{\eta}(z;\theta)]-\mathbb{E}_{z\sim\mathsf{P}^{o}_{n}}[l_{\eta}(z;\theta)]\rvert$, where $l_{\eta}(z;\theta)=\inf_{z\in\mathcal{Z}}[\eta d^{p}(z,z^{\prime})-l(z;\theta)]$ is called the Moreau-Yosida regularization of $-l$ with parameter $1/\eta$. We show that, for all $\eta\geq 0$, all $l_{\eta}$ are uniformly bounded. We then use Hoeffding’s inequality to obtain concentration. Detailed proof is in Section B.2.
Next, when Assumption 2 is in place, we can show that $g(\theta)\coloneqq\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex w.r.t. the positive definite norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. Further, by the property of supremum, we can show that $\mathcal{L}^{\mathrm{W}}$ is $\gamma\lambda$-strongly convex but w.r.t. $\lVert\cdot\rVert_{2}$. A detailed proof is provided in Section B.3.
Decompose $\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}}_{n})-\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}})$ into three terms: $\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}}_{n};\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}}_{n};\rho)$, $\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}}_{n};\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}};\rho)$, and $\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}};\rho)-\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}};\rho)$. The second term is non-positive since $\theta^{\mathrm{W}}_{n}$ is the minimizer of $\mathcal{L}^{\mathrm{W}}_{n}$. Now we apply the concentration of the WDPO loss function (see Lemma 9 in Section B.2) to $\lvert\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}}_{n};\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}}_{n};\rho)\rvert$ and $\lvert\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}};\rho)-\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}};\rho)\rvert$. Finally, we use the property of strongly convex function (Lemma 5) on $\mathcal{L}^{\mathrm{W}}$ to acquire the policy parameter convergence. The detailed proof is in Section B.4. ∎
@@
A matching result can be derived as a special case of Chowdhury et al., (2024, Theorem 4.2). We provide an independent proof with precise constants in Section B.1.
@@
We would like to note that the estimation error rate of convergence for WDPO is $\lVert\theta^{\mathrm{W}}_{n}-\theta^{\mathrm{W}}\rVert_{2}=O(n^{-1/4})$, from Theorem 1. The estimation error rate of convergence for (non-robust) DPO is $\lVert\theta_{n}-\theta^{*}\rVert_{\Sigma_{\mathcal{D}}+\lambda I}=O(n^{-1/2})$, from Proposition 1. So, the estimation error rate of convergence for WDPO is worse than that of (non-robust) DPO. This arises due to significant challenges exclusive to the robust setting. For example, for the non-robust DPO, we can calculate the closed-form expression of $\nabla_{\theta}(1/n)\sum_{i=1}^{n}l(z_{i};\theta)$ (see Eq. 20). This allows us to write $\lVert\nabla_{\theta}(1/n)\sum_{i=1}^{n}l(z_{i};\theta^{*})\rVert_{(\Sigma_{\mathcal{D}}+\lambda I)^{-1}}$ in quadratic form and then obtain a concentration using Bernstein’s inequality. However, for WDPO, we note that $\nabla_{\theta}\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}})\neq\sup_{\mathsf{P}\in\mathcal{P}_{\mathsf{W}_{p}}}\nabla_{\theta}\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta^{\mathrm{W}})]$, and the non-robust approach will not work for the robust setting. Developing analysis techniques to achieve a better rate of convergence for robust DPO is an open question.
@@ 5.2 Estimation Error for KLDPO
Let $\theta^{\mathrm{KL}}\in\operatorname*{argmin}_{\theta\in\Theta}\mathcal{L}^{\mathrm{KL}}(\theta;\rho)$, and let its empirical counterpart be $\theta^{\mathrm{KL}}_{n}\in\operatorname*{argmin}_{\theta\in\Theta}\mathcal{L}^{\mathrm{KL}}_{n}(\theta;\rho)$. The convergence analysis for the KLDPO loss and policy parameter closely parallels that of Wasserstein DPO. We present the main theorems below and defer detailed proofs to Appendix C.
@@
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$. $\lambda$ is the regularity condition number defined in Assumption 2, $0<\lambda\leq\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}^{o}})$. $\underline{\lambda},\overline{\lambda}$ are some universal constants, and $L$ is an upper bound on the loss function $l$.
@@
The exponential constant in the upper bound is a characteristic of distributional robust optimization with KL uncertainty set Hu and Hong, (2013, Proposition 2). Similar exponential constants appear in the theoretical analysis of the distributionally robust RL (Zhou et al.,, 2021; Yang et al.,, 2022; Panaganti and Kalathil,, 2022; Xu et al.,, 2023). Both WDPO and KLDPO have $O(n^{-1/4})$ policy parameter convergence. An empirical comparison is given in Section 7.
@@ 6 Tractable (Approximate) Algorithms
While our distributionally robust DPO formulations enjoy finite-sample guarantees, it is computationally challenging to solve the min-max objective of Eq. 6 using stochastic gradient descent methods. Though many min-max optimization problems can be solved by alternating gradient descent methods, our problem is not directly amenable to such an approach as we do not have direct control over the data distribution $\mathsf{P}\in\mathcal{P}(\rho;\mathsf{P}^{o})$ which is not parameterized. Moreover, the preference data are generated according to the nominal distribution $\mathsf{P}^{o}$ ,and we do not have data samples from any other distributions in the uncertainty set $\mathcal{P}(\rho;\mathsf{P}^{o})$. To overcome this challenge, we introduce principled tractable algorithms to solve WDPO and KLDPO.
Tractable WDPO: The connection between Wasserstein distributionally robust optimization (DRO) and regularization has been established in various settings by many (Mohajerin Esfahani and Kuhn,, 2018; Shafieezadeh-Abadeh et al.,, 2019; Chen and Paschalidis,, 2018). We leverage the recent progress in Wasserstein theory on connecting Wasserstein distributionally robust optimization to regularization. For $p$-Wasserstein DRO, $p\in(1,\infty]$, Gao et al., (2022) shows that for a broad class of loss functions, possibly non-convex and non-smooth, with high probability, the Wasserstein DRO is asymptotically equivalent to variation regularization. In particular, an immediate consequence of Gao et al., (2022, Theorem 1) is that, when $p=2$,
where $\rho_{n}=O(1/\sqrt{n})$. That is, one can solve the Wasserstein DRO objective by adding a gradient regularization to the empirical risk minimization (ERM) loss, $\mathbb{E}_{z\sim\mathsf{P}^{o}_{n}}[l(z;\theta)]$. Based on this, we propose a tractable WDPO algorithm in Algorithm 1.
@@
3:     Calculate the non-robust DPO loss $\mathcal{L}^{\mathrm{DPO}}(\pi_{\theta};\mathcal{D})$ according to Eq. 3
@@
$\mathcal{R}(\pi_{\theta};\mathcal{D})=\rho_{o}(\mathbb{E}_{z\sim\mathcal{D}}\lVert\nabla_{z}l(z;\theta)\rVert_{2}^{2})^{1/2}$
@@
6.1 Tractable KLDPO: Approximate Dual Solution
Tractable KLDPO: The following proposition shows that we can approximate the worst-case probability distribution in a KL uncertainty set w.r.t. a given loss function. Similar results can also be found in distributionally robust reinforcement learning literature (e.g., Gadot et al., (2024)).
@@ 6.1 Tractable KLDPO: Approximate Dual Solution
We defer the formal proof of Proposition 2 to Appendix D. It can be viewed as a re-weighting threshold: extreme losses are more biased towards the baseline empirical DPO loss. $\tau$ controls the intensity of re-weighting, acting as a temperature parameter. Based on Proposition 2, we propose a tractable KLDPO algorithm in Algorithm 2.
Algorithm 2 KLDPO Algorithm
2:while $\theta$ has not converged do
3:     Approximate the worst-case kernel
$\underline{\mathsf{P}}(i)\propto\mathrm{exp}\left((1/\tau)(l(z_{i};\theta)-(1/n)\textstyle\sum\nolimits_{i=1}^{n}l(z_{i};\theta))\right)$
4:     Calculate the approximate KLDPO loss
6:Output: $\pi_{\theta}$
@@ 7 Experiments
We conduct experiments across three distinct settings that vary in dataset scale, model size, and the degree of distribution shift. For example, we fine-tune LLaMA-3.2-1B-Instruct, LLaMA-3.2-3B-Instruct, and LLaMA-3.1-8B-Instruct models on prompts from the HelpSteer2 dataset (Wang et al., 2024b, ), using preferences derived from the ArmoRM reward model (Wang et al., 2024a, ), and evaluate them on the OpenLLM Leaderboard v2 (Fourrier et al.,, 2024). Additional evaluations are provided in Appendix E. We provide the code at https://github.com/TheBlackCat22/distributionally_robust_dpo.
@@ 7.1 Experimental Setup
Emotion Alignment: We use the Emotion dataset (Saravia et al.,, 2018) to train a GPT-2 model (Radford et al.,, 2019) with a classification head for multi-label classification over five emotions: sadness, joy, love, anger, fear. The resulting sigmoid outputs are used as a multi-objective reward model for the remainder of this experiment. We also take a GPT-2 model and perform supervised fine-tuning (SFT) with the Emotion dataset to obtain our base model for preference alignment. To construct preference data, we mix objectives derived from our reward model. Specifically, we consider two reward objectives, $r_{1},r_{2}$ and define two mixture reward functions (1) convex mixing $r^{*}_{\mathrm{convex}}(\alpha)\coloneqq\alpha\cdot r_{1}+(1-\alpha)\cdot r_{2}$ and (2) geometric mixing $r^{*}_{\mathrm{geometric}}(\alpha)\coloneqq r_{1}^{\alpha}\cdot r_{2}^{1-\alpha}$. For both reward functions, we generate two completions per prompt and assign preference labels using the Bradley-Terry (BT) model parameterized by $r^{*}(\alpha^{o})$ for a chosen $\alpha^{o}\in[0,1]$.
ArmoRM Multi-objective Alignment: We use the Absolute-Rating Multi-Objective Reward Model (ArmoRM) (Wang et al., 2024a, ) to define reward preferences, selecting pairs of equally weighted objectives (e.g., honesty, verbosity, safety) from its 19-dimensional first-stage outputs. Using Meta LLaMA-3.2-1B-Instruct as the base model, we generate two completions per prompt from the HelpSteer2 dataset (Wang et al., 2024b, ) and train models on preferences derived from the convex mixing of these reward pairs. We evaluate all models on five individual ArmoRM objectives, three of which are unseen during training, to simulate preference shift.
Leaderboard Alignment: We fine-tune LLaMA-3.2-1B-Instruct, LLaMA-3.2-3B-Instruct, and LLaMA-3.1-8B-Instruct models using preference data derived from the scalar rewards produced by the second stage of the ArmoRM reward model (Wang et al., 2024a, ). For each prompt from the HelpSteer2 dataset, we generate 10 responses, score them with ArmoRM, and constructe preference pairs by selecting the highest- and lowest-scoring completions. The models are evaluated on the OpenLLM Leaderboard v2 (Fourrier et al.,, 2024) using the LM Evaluation Harness (Gao et al.,, 2024).

### v3

@@ 5.1 Estimation Error for WDPO
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$ and $K=\lvert\log\sigma(-4\beta B)\rvert$, $\lambda$ is the regularity number defined in Assumption˜2.
@@
Strong duality of Wasserstein DRO (see Gao and Kleywegt (2022) and Corollary˜1) helps us reduce the difference $\left\lvert\mathcal{L}^{\mathrm{W}}(\theta;\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta;\rho)\right\rvert$ to the concentration $\lvert\mathbb{E}_{z\sim\mathsf{P}^{o}}[l_{\eta}(z;\theta)]-\mathbb{E}_{z\sim\mathsf{P}^{o}_{n}}[l_{\eta}(z;\theta)]\rvert$, where $l_{\eta}(z;\theta)=\inf_{z\in\mathcal{Z}}[\eta d^{p}(z,z^{\prime})-l(z;\theta)]$ is called the Moreau-Yosida regularization of l-l with parameter $1/\eta$. We show that, for all $\eta\geq 0$, all $l_{\eta}$ are uniformly bounded. We then use Hoeffding’s inequality to obtain concentration. Detailed proof is in Section˜B.2.
Next, when Assumption˜2 is in place, we can show that $g(\theta)\coloneqq\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex w.r.t. the positive definite norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. Further, by the property of supremum, we can show that $\mathcal{L}^{\mathrm{W}}$ is $\gamma\lambda$-strongly convex but w.r.t. $\lVert\cdot\rVert_{2}$. A detailed proof is provided in Section˜B.3.
Decompose $\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}}_{n})-\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}})$ into three terms: $\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}}_{n};\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}}_{n};\rho)$, $\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}}_{n};\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}};\rho)$, and $\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}};\rho)-\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}};\rho)$. The second term is non-positive since $\theta^{\mathrm{W}}_{n}$ is the minimizer of $\mathcal{L}^{\mathrm{W}}_{n}$. Now we apply the concentration of the WDPO loss function (see Lemma˜9 in Section˜B.2) to $\lvert\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}}_{n};\rho)-\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}}_{n};\rho)\rvert$ and $\lvert\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}};\rho)-\mathcal{L}^{\mathrm{W}}(\theta^{\mathrm{W}};\rho)\rvert$. Finally, we use the property of strongly convex function (Lemma˜5) on $\mathcal{L}^{\mathrm{W}}$ to acquire the policy parameter convergence. The detailed proof is in Section˜B.4. ∎
@@
A matching result can be derived as a special case of Chowdhury et al. (2024, Theorem 4.2). We provide an independent proof with precise constants in Section˜B.1.
@@
We would like to note that the estimation error rate of convergence for WDPO is $\lVert\theta^{\mathrm{W}}_{n}-\theta^{\mathrm{W}}\rVert_{2}=O(n^{-1/4})$, from Theorem˜1. The estimation error rate of convergence for (non-robust) DPO is $\lVert\theta_{n}-\theta^{*}\rVert_{\Sigma_{\mathcal{D}}+\lambda I}=O(n^{-1/2})$, from Proposition˜1. So, the estimation error rate of convergence for WDPO is worse than that of (non-robust) DPO. This arises due to significant challenges exclusive to the robust setting. For example, for the non-robust DPO, we can calculate the closed-form expression of $\nabla_{\theta}(1/n)\sum_{i=1}^{n}l(z_{i};\theta)$ (see Eq.˜20). This allows us to write $\lVert\nabla_{\theta}(1/n)\sum_{i=1}^{n}l(z_{i};\theta^{*})\rVert_{(\Sigma_{\mathcal{D}}+\lambda I)^{-1}}$ in quadratic form and then obtain a concentration using Bernstein’s inequality. However, for WDPO, we note that $\nabla_{\theta}\mathcal{L}^{\mathrm{W}}_{n}(\theta^{\mathrm{W}})\neq\sup_{\mathsf{P}\in\mathcal{P}_{\mathsf{W}_{p}}}\nabla_{\theta}\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta^{\mathrm{W}})]$, and the non-robust approach will not work for the robust setting. Developing analysis techniques to achieve a better rate of convergence for robust DPO is an open question.
@@ 5.2 Estimation Error for KLDPO
Let $\theta^{\mathrm{KL}}\in\operatorname*{argmin}_{\theta\in\Theta}\mathcal{L}^{\mathrm{KL}}(\theta;\rho)$, and let its empirical counterpart be $\theta^{\mathrm{KL}}_{n}\in\operatorname*{argmin}_{\theta\in\Theta}\mathcal{L}^{\mathrm{KL}}_{n}(\theta;\rho)$. The convergence analysis for the KLDPO loss and policy parameter closely parallels that of Wasserstein DPO. We present the main theorems below and defer detailed proofs to Appendix˜C.
@@
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$. $\lambda$ is the regularity condition number defined in Assumption˜2, $0<\lambda\leq\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}^{o}})$. $\underline{\lambda},\overline{\lambda}$ are some universal constants, and $L$ is an upper bound on the loss function $l$.
@@
The exponential constant in the upper bound is a characteristic of distributional robust optimization with KL uncertainty set Hu and Hong (2013, Proposition 2). Similar exponential constants appear in the theoretical analysis of the distributionally robust RL (Zhou et al., 2021; Yang et al., 2022; Panaganti and Kalathil, 2022; Xu et al., 2023). Both WDPO and KLDPO have $O(n^{-1/4})$ policy parameter convergence. An empirical comparison is given in Section˜7.
@@ 6 Tractable (Approximate) Algorithms
3:  Calculate the non-robust DPO loss $\mathcal{L}^{\mathrm{DPO}}(\pi_{\theta};\mathcal{D})$ according to Eq.˜3
@@
$\mathcal{R}(\pi_{\theta};\mathcal{D})=\rho_{o}(\mathbb{E}_{z\sim\mathcal{D}}\lVert\nabla_{z}l(z;\theta)\rVert_{2}^{2})^{1/2}\vskip-2.84544pt$
@@
Algorithm 2 KLDPO Algorithm
2:while $\theta$ has not converged do
3:  Approximate the worst-case kernel
$\underline{\mathsf{P}}(i)\propto\mathrm{exp}((1/\tau)($ $l(z_{i};\theta)$
$-(1/n)\textstyle\sum\nolimits_{i=1}^{n}l(z_{i};\theta)))$
4:  Calculate the approximate KLDPO loss
6:Output: $\pi_{\theta}$
While our distributionally robust DPO formulations enjoy finite-sample guarantees, it is computationally challenging to solve the min-max objective of Eq.˜6 using stochastic gradient descent methods. Though many min-max optimization problems can be solved by alternating gradient descent methods, our problem is not directly amenable to such an approach as we do not have direct control over the data distribution $\mathsf{P}\in\mathcal{P}(\rho;\mathsf{P}^{o})$ which is not parameterized. Moreover, the preference data are generated according to the nominal distribution $\mathsf{P}^{o}$ ,and we do not have data samples from any other distributions in the uncertainty set $\mathcal{P}(\rho;\mathsf{P}^{o})$. To overcome this challenge, we introduce principled tractable algorithms to solve WDPO and KLDPO.
Tractable WDPO: The connection between Wasserstein distributionally robust optimization (DRO) and regularization has been established in various settings by many (Mohajerin Esfahani and Kuhn, 2018; Shafieezadeh-Abadeh et al., 2019; Chen and Paschalidis, 2018). We leverage the recent progress in Wasserstein theory on connecting Wasserstein distributionally robust optimization to regularization. For $p$-Wasserstein DRO, $p\in(1,\infty]$, Gao et al. (2022) shows that for a broad class of loss functions, possibly non-convex and non-smooth, with high probability, the Wasserstein DRO is asymptotically equivalent to variation regularization. In particular, an immediate consequence of Gao et al. (2022, Theorem 1) is that, when $p=2$,
where $\rho_{n}=O(1/\sqrt{n})$. That is, one can solve the Wasserstein DRO objective by adding a gradient regularization to the empirical risk minimization (ERM) loss, $\mathbb{E}_{z\sim\mathsf{P}^{o}_{n}}[l(z;\theta)]$. Based on this, we propose a tractable WDPO algorithm in Algorithm˜1.
Tractable KLDPO: The following proposition shows that we can approximate the worst-case probability distribution in a KL uncertainty set w.r.t. a given loss function. Similar results can also be found in distributionally robust reinforcement learning literature (e.g., Gadot et al. (2024)).
@@
We defer the formal proof of Proposition˜2 to Appendix˜D. It can be viewed as a re-weighting threshold: extreme losses are more biased towards the baseline empirical DPO loss. $\tau$ controls the intensity of re-weighting, acting as a temperature parameter. Based on Proposition˜2, we propose a tractable KLDPO algorithm in Algorithm˜2.
@@ 7 Experiments
Figure 2: DPO, WDPO, and KLDPO in Emotion Alignment. Models are trained on preferences derived from convex (left two plots) and geometric (right two plots) mixtures of anger and fear objectives from the Emotion dataset (Saravia et al., 2018). To simulate preference shift, evaluation is performed at mixing coefficients $\alpha\neq\alpha_{o}$, where $\alpha_{o}=0.1$ is used during training. See Section˜7.1.
We conduct experiments across three distinct settings that vary in dataset scale, model size, and the degree of distribution shift. For example, we fine-tune LLaMA-3.2-1B-Instruct, LLaMA-3.2-3B-Instruct, and LLaMA-3.1-8B-Instruct models on prompts from the HelpSteer2 dataset (Wang et al., 2024b), using preferences derived from the ArmoRM reward model (Wang et al., 2024a), and evaluate them on the OpenLLM Leaderboard v2 (Fourrier et al., 2024). Additional evaluations are provided in Appendix˜E. We provide the code at https://github.com/TheBlackCat22/distributionally_robust_dpo.
@@ 7.1 Experimental Setup
Emotion Alignment: We use the Emotion dataset (Saravia et al., 2018) to train a GPT-2 model (Radford et al., 2019) with a classification head for multi-label classification over five emotions: sadness, joy, love, anger, fear. The resulting sigmoid outputs are used as a multi-objective reward model for the remainder of this experiment. We also take a GPT-2 model and perform supervised fine-tuning (SFT) with the Emotion dataset to obtain our base model for preference alignment. To construct preference data, we mix objectives derived from our reward model. Specifically, we consider two reward objectives, $r_{1},r_{2}$ and define two mixture reward functions (1) convex mixing $r^{*}_{\mathrm{convex}}(\alpha)\coloneqq\alpha\cdot r_{1}+(1-\alpha)\cdot r_{2}$ and (2) geometric mixing $r^{*}_{\mathrm{geometric}}(\alpha)\coloneqq r_{1}^{\alpha}\cdot r_{2}^{1-\alpha}$. For both reward functions, we generate two completions per prompt and assign preference labels using the Bradley-Terry (BT) model parameterized by $r^{*}(\alpha^{o})$ for a chosen $\alpha^{o}\in[0,1]$.
ArmoRM Multi-objective Alignment: We use the Absolute-Rating Multi-Objective Reward Model (ArmoRM) (Wang et al., 2024a) to define reward preferences, selecting pairs of equally weighted objectives (e.g., honesty, verbosity, safety) from its 19-dimensional first-stage outputs. Using Meta LLaMA-3.2-1B-Instruct as the base model, we generate two completions per prompt from the HelpSteer2 dataset (Wang et al., 2024b) and train models on preferences derived from the convex mixing of these reward pairs. We evaluate all models on five individual ArmoRM objectives, three of which are unseen during training, to simulate preference shift.
Leaderboard Alignment: We fine-tune LLaMA-3.2-1B-Instruct, LLaMA-3.2-3B-Instruct, and LLaMA-3.1-8B-Instruct models using preference data derived from the scalar rewards produced by the second stage of the ArmoRM reward model (Wang et al., 2024a). For each prompt from the HelpSteer2 dataset, we generate 10 responses, score them with ArmoRM, and constructe preference pairs by selecting the highest- and lowest-scoring completions. The models are evaluated on the OpenLLM Leaderboard v2 (Fourrier et al., 2024) using the LM Evaluation Harness (Gao et al., 2024).